OpenStudy (anonymous):

Question 1 Factor x2 – 21x – 100 a. prime b. (x – 50)(x + 2) c. (x + 4)(x – 25) d. (x + 25)(x – 4) QUESTION 2 Factor x2 + 5x + 6. a.(x + 2)(x + 3) b. (x + 1)(x + 6) c. (x + 1)(x + 5) d. prime QUESTION 3 Factor x2 + 7x + 12. a. (x + 3)(x + 4) b. (x + 1)(x + 7) c. prime d. (x + 2)(x + 6) QUESTION 4 Factor 6x2 – 7x – 4. a. (2x + 2)(3x – 2) b. (6x + 2)(x – 2) c. The polynomial is prime. d. (6x – 4)(x + 1) QUESTION 5 Factor 2x2 + 5x + 3. a. (2x + 1)(x + 3) b. (2x + 3)(x + 1) c. The polynomial is prime. d. (2x + 5)(x + 1)

OpenStudy (anonymous):

It's probably easiest just to multiply the MC choice out until you find the solution. BUT that won't help you learn to factor.

OpenStudy (anonymous):

If you multiply and keep track of there the term come from, then you know what to look for when factoring. For example with \[(x+1)(x-2) = x^2-x-2\] the x^2 came from the first terms in the brackets x*x = x^2, the (-1) part of -x came from adding the last parts of the brackets, and finally the -2 came from (+1)*(-2) multiplying that last terms in the brackets.